# Counterfactual Simulatability of LLM Explanations for Generation Tasks

### v2

@@ 1 Introduction
While large language models (LLMs) have proven effective for a diverse range of applications, their outputs still often contain hallucinations of unsupported information (Ji et al., 2023) or biases (Sheng et al., 2021) that hinder their reliability on critical language generation tasks. At the same time, it is important that users themselves can accurately evaluate the model capabilities, knowledge, and reliability (Steyvers et al., 2025). Particularly in high stakes domains, such as medical applications, misunderstandings or the lack of ability to predict model behavior on unseen inputs can pose disastrous risks to users (Michalowski et al., 2024).
To anticipate these risks, recent work has turned attention to evaluating the reliability of LLM explanations (Madsen et al., 2024; Turpin et al., 2023; Kunz and Kuhlmann, 2024). In particular, Chen et al. (2024) evaluates the counterfactual simulatability of LLM explanations for yes/no question answering. Counterfactual simulatability is a measure of how well a model’s explanation allows humans to correctly infer the model’s predictions on simulatable counterfactuals (unseen inputs where the explanation should enable the user to confidently guess the model’s output). Furthermore, according to Chen et al. (2024), the counterfactual simulatability of an explanation can be decomposed into simulation generality, a measure of the diversity of simulatable counterfactuals and simulation precision, a measure of the proportion of these counterfactuals for which humans correctly infer the model’s output. Ideal model explanations should balance both generality and precision.
@@
We apply the framework to two tasks: news summarization using CNN/DM (Nallapati et al., 2016) and medical suggestion generation using the Taiwan e-Hospital Dataset (Chen et al., 2022b). These tasks involve different trade-offs between generality and precision. Explanations for news summarization can be highly general as models may employ similar approaches for documents that contain common elements (e.g., dates, quotes) but differ in content. In contrast, medical suggestion explanations may be less general as the model’s response can be highly dependent on the user input (i.e., minor variances in expressed symptoms can lead to different suggestions). This can lead to a more limited set of simulatable counterfactuals, though potentially being more precise in its outputs. With these generation tasks, we conduct an initial evaluation of LLMs’ explanations and assess where models’ explanations in these generation tasks fail.
@@ 2 Related Work
Human mental models. Humans form mental models of the physical world as a whole (Gentner and Stevens, 2014) as well as specific technologies (e.g., Payne, 1991; Du et al., 2018; Lei et al., 2016) through their past experiences and observations. Specifically with regard to artificial intelligence systems, explanations of model predictions have been considered high quality if they provide users with an accurate and generalizable understanding of the system in the form of mental models (Rutjes et al., 2019; Merry et al., 2021). When such explanations are successful, they can improve users’ ability to effectively use AI models (Vasconcelos et al., 2022; Senoner et al., 2024) as well as to anticipate and correct undesirable model behaviors, such as biases and incorrect predictions (Bansal et al., 2019). We specifically evaluate LLMs’ explanations in generation tasks considering their ability to help form mental models.
Explanation evaluation. A variety of different approaches have been used to evaluate the quality and utility of model-generated natural language explanations and rationales. Studies evaluating natural language explanations similar to word attributions (Huang et al., 2023; Madsen et al., 2024) as well as unconstrained explanations (Turpin et al., 2023) have focused on faithfulness measures. In particular, work has proposed metrics for dimensions including comprehensiveness (DeYoung et al., 2020), sufficiency (DeYoung et al., 2020), alignment with human rationales (Fayyaz et al., 2024), plausibility (Wojciechowski et al., 2024), and scrutability (Xu et al., 2023) among others. In contrast to intrinsic measures of model explanations, other work evaluates explanations extrinsically through impact on task performance (e.g., Camburu et al., 2018; Wei et al., 2022; Krishna et al., 2023). Among these topics, relatively few works have investigated natural language explanations in language generation tasks; such studies include dialogue responses (Zhou et al., 2021), dialogue understanding (Gao et al., 2024), and more prominently, open-ended question answering (Ho et al., 2023; Fragkathoulas and Chlapanis, 2024; Lyu et al., 2023). While Chen et al. (2024) introduces counterfactual simulatability in a binary classification setting, we propose a novel framework for counterfactual simulatability in language generation settings.
@@ 3 Counterfactual Simulatability for Generation Tasks
Explanations can be evaluated by considering whether an observer, having seen a model’s explanation for some input, can infer (i.e., simulate) the model’s output on a related counterfactual input. For instance, if a user asks “do dolphins swim?” and a model answers “yes” with the explanation “all aquatic animals swim”, then the user would infer that when asked the counterfactual “do starfish swim?”, the model will similarly answer “yes”. If, in reality, the model answers “no” to this question, then, as Chen et al. (2024) notes, the explanation is ineffective because it creates a mental model that is inconsistent with the model’s behavior, even though the answer may be factually correct. More specifically, counterfactual simulatability measures how accurate these mental models are on simulatable counterfactuals, unseen inputs where the explanation allows the user to confidently predict the model’s output (Chen et al., 2024).
@@ 4.1 Datasets
We select two generative tasks that represent practical LLM use cases and whose natural language explanations differ in generality and precision. For the first task of news summarization, we use the CNN/DM dataset (Nallapati et al., 2016). Summarization is chosen as it is a well-studied application for language generation. Additionally, this task allows us to study high-level, abstract explanations since summarization does not rely on facts learned during the LLM’s pre-training, rather its ability to extract key elements from the input document.
Conversely, we also investigate medical suggestion, a domain where explanations are critical. For this task, we use the Taiwan e-Hospital Dataset, a collection of 86,399 Mandarin question answer pairs from the online health website Taiwan e-Hospital (Chen et al., 2022b). The data was translated into English using the Google Translate API. We select this specific dataset because each sample consists of a question, suggestion, explanation triplet, ensuring that the questions are complex enough such that explanations are helpful. Furthermore, the dataset also mirrors the practical setting we aim to study, where everyday users interact with LLMs to seek general medical advice without any specialized knowledge.
@@ 4.3 Explanations
Through prompting (instructions, few-shot examples), we guide the LLMs to generate explanations that emphasize the difference in generality and precision across tasks. For news summarization, to be highly generalizable, we encourage the model to explain its decision process at a high-level, identifying abstract elements while avoiding reference to specific topics. For medical suggestion, we instruct the model to first identify important details in the user’s question, propose possible underlying causes, and suggest recommended actions accordingly. These approaches reflect the skill-based vs. knowledge-based nature of our tasks. Additionally, following Chen et al. (2024), we experiment with both Chain-of-thought and Post-hoc prompting to produce explanations (Camburu et al., 2018). The prompts used are provided in Appendix D.
@@ 4.4 Estimating mental models
Generation tasks are challenging for counterfactual simulatability because mental models are extremely difficult to estimate given the large space of possible outputs. Motivated by previous work on atomic units (Wright et al., 2022; Chen et al., 2022a; Kamoi et al., 2023), we introduce a method for estimating mental models by decomposing explanations into units of information as a proxy.

### v3

@@ 1 Introduction
While large language models (LLMs) have proven effective for a diverse range of applications, their outputs still often contain hallucinations of unsupported information (ji2023survey) or biases (sheng-etal-2021-societal) that hinder their reliability on critical language generation tasks. At the same time, it is important that users themselves can accurately evaluate the model capabilities, knowledge, and reliability (steyvers-think). Particularly in high stakes domains, such as medical applications, misunderstandings or the lack of ability to predict model behavior on unseen inputs can pose disastrous risks to users (michalowski-explain).
To anticipate these risks, recent work has turned attention to evaluating the reliability of LLM explanations (madsen-etal-2024-self; turpin-cot; kunz-kuhlmann-2024-properties). In particular, chen-simulatbility evaluates the counterfactual simulatability of LLM explanations for yes/no question answering. Counterfactual simulatability is a measure of how well a model’s explanation allows humans to correctly infer the model’s predictions on simulatable counterfactuals (unseen inputs where the explanation should enable the user to confidently guess the model’s output). Furthermore, according to chen-simulatbility, the counterfactual simulatability of an explanation can be decomposed into simulation generality, a measure of the diversity of simulatable counterfactuals and simulation precision, a measure of the proportion of these counterfactuals for which humans correctly infer the model’s output. Ideal model explanations should balance both generality and precision.
@@
We apply the framework to two tasks: news summarization using CNN/DM (nallapati-etal-2016-abstractive) and medical suggestion generation using the Taiwan e-Hospital Dataset (chen-etal-2022-learning-generate). These tasks involve different trade-offs between generality and precision. Explanations for news summarization can be highly general as models may employ similar approaches for documents that contain common elements (e.g., dates, quotes) but differ in content. In contrast, medical suggestion explanations may be less general as the model’s response can be highly dependent on the user input (i.e., minor variances in expressed symptoms can lead to different suggestions). This can lead to a more limited set of simulatable counterfactuals, though potentially being more precise in its outputs. With these generation tasks, we conduct an initial evaluation of LLMs’ explanations and assess where models’ explanations in these generation tasks fail.
@@ 2 Related Work
Human mental models. Humans form mental models of the physical world as a whole (gentner-mental) as well as specific technologies (e.g., payne-mental; du-voice; lei-mobile) through their past experiences and observations. Specifically with regard to artificial intelligence systems, explanations of model predictions have been considered high quality if they provide users with an accurate and generalizable understanding of the system in the form of mental models (rutjes-xai; merry-xai). When such explanations are successful, they can improve users’ ability to effectively use AI models (vasconcelos-cost; senoner-collab) as well as to anticipate and correct undesirable model behaviors, such as biases and incorrect predictions (bansal-accuracy). We specifically evaluate LLMs’ explanations in generation tasks considering their ability to help form mental models.
Explanation evaluation. A variety of different approaches have been used to evaluate the quality and utility of model-generated natural language explanations and rationales. Studies evaluating natural language explanations similar to word attributions (huang-explain; madsen-etal-2024-self) as well as unconstrained explanations (turpin-cot) have focused on faithfulness measures. In particular, work has proposed metrics for dimensions including comprehensiveness (deyoung-eraser), sufficiency (deyoung-eraser), alignment with human rationales (fayyaz-alignment), plausibility (wojciechowski2024faithful), and scrutability (zhichao-scrutability) among others. In contrast to intrinsic measures of model explanations, other work evaluates explanations extrinsically through impact on task performance (e.g., camburu-esnli; wei-cot; krishna-posthoc). Among these topics, relatively few works have investigated natural language explanations in language generation tasks; such studies include dialogue responses (zhou-commonsense), dialogue understanding (gao-dialogue), and more prominently, open-ended question answering (ho-wikiwhy; fragkathoulas-qa; lyu-faithful). While chen-simulatbility introduces counterfactual simulatability in a binary classification setting, we propose a novel framework for counterfactual simulatability in language generation settings.
@@ 3 Counterfactual Simulatability for Generation Tasks
Explanations can be evaluated by considering whether an observer, having seen a model’s explanation for some input, can infer (i.e., simulate) the model’s output on a related counterfactual input. For instance, if a user asks “do dolphins swim?” and a model answers “yes” with the explanation “all aquatic animals swim”, then the user would infer that when asked the counterfactual “do starfish swim?”, the model will similarly answer “yes”. If, in reality, the model answers “no” to this question, then, as chen-simulatbility notes, the explanation is ineffective because it creates a mental model that is inconsistent with the model’s behavior, even though the answer may be factually correct. More specifically, counterfactual simulatability measures how accurate these mental models are on simulatable counterfactuals, unseen inputs where the explanation allows the user to confidently predict the model’s output (chen-simulatbility).
@@ 4.1 Datasets
We select two generative tasks that represent practical LLM use cases and whose natural language explanations differ in generality and precision. For the first task of news summarization, we use the CNN/DM dataset (nallapati-etal-2016-abstractive). Summarization is chosen as it is a well-studied application for language generation. Additionally, this task allows us to study high-level, abstract explanations since summarization does not rely on facts learned during the LLM’s pre-training, rather its ability to extract key elements from the input document.
Conversely, we also investigate medical suggestion, a domain where explanations are critical. For this task, we use the Taiwan e-Hospital Dataset, a collection of 86,399 Mandarin question answer pairs from the online health website Taiwan e-Hospital (chen-etal-2022-learning-generate). The data was translated into English using the Google Translate API. We select this specific dataset because each sample consists of a question, suggestion, explanation triplet, ensuring that the questions are complex enough such that explanations are helpful. Furthermore, the dataset also mirrors the practical setting we aim to study, where everyday users interact with LLMs to seek general medical advice without any specialized knowledge.
@@ 4.3 Explanations
Through prompting (instructions, few-shot examples), we guide the LLMs to generate explanations that emphasize the difference in generality and precision across tasks. For news summarization, to be highly generalizable, we encourage the model to explain its decision process at a high-level, identifying abstract elements while avoiding reference to specific topics. For medical suggestion, we instruct the model to first identify important details in the user’s question, propose possible underlying causes, and suggest recommended actions accordingly. These approaches reflect the skill-based vs. knowledge-based nature of our tasks. Additionally, following chen-simulatbility, we experiment with both Chain-of-thought and Post-hoc prompting to produce explanations (camburu-esnli). The prompts used are provided in Appendix D.
@@ 4.4 Estimating mental models
Generation tasks are challenging for counterfactual simulatability because mental models are extremely difficult to estimate given the large space of possible outputs. Motivated by previous work on atomic units (wright2022generating; chen2022generating; kamoi2023wice), we introduce a method for estimating mental models by decomposing explanations into units of information as a proxy.